sunisround:

limit of (((x-5)^2)-25))/(x) as x approaches 0. I need the steps on how to solve it.

sillybilly123:

\(\lim\limits_{x \to 0} \dfrac{(x-5)^2-25}{x }\) \(= \lim\limits_{x \to 0} \dfrac{(x^2 - 10 x + 25)-25}{x }\) \(= \lim\limits_{x \to 0} \dfrac{x(x - 10 )}{x }\) \(= \lim\limits_{x \to 0} (x - 10 ) = - 10\)